### v1

# Reinforcement Learning Finetunes Small Subnetworks in Large Language Models

Abstract
Reinforcement learning (RL) yields substantial improvements in large language models’ (LLMs) downstream task performance and alignment with human values. Surprisingly, such large gains result from updating only a small subnetwork comprising just 5%-30% of the parameters, with the rest effectively unchanged. We refer to this phenomenon as parameter update sparsity induced by RL. It is observed across all 7 widely-used RL algorithms (e.g., PPO, GRPO, DPO) and all 10 LLMs from different families in our experiments. This sparsity is intrinsic and occurs without any explicit sparsity-promoting regularizations or architectural constraints. Finetuning the subnetwork alone recovers the test accuracy, and, remarkably, produces a model nearly identical to the one obtained via full finetuning. The subnetworks from different random seeds, training data, and even RL algorithms show substantially greater overlap than expected by chance. Our analysis suggests that this sparsity is not due to updating only a subset of layers; instead, nearly all parameter matrices receive similarly sparse updates. Moreover, the updates to almost all parameter matrices are nearly full-rank, suggesting RL updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent. We conjecture that the this update sparsity can be primarily attributed to training on data that is near the policy distribution; techniques that encourage the policy to remain close to the pretrained model, such as the KL regularization and gradient clipping, have limited impact.
1 Introduction
Reinforcement learning (RL) (Sutton et al., 1998; Ouyang et al., 2022; Ziegler et al., 2020; Ramamurthy et al., 2023; Sun et al., 2024; Zhou et al., 2025) is an important post-pretraining stage for adapting large language models (LLMs) to solving complex reasoning problems (Lightman et al., 2023; Wang et al., 2025a, c; Cui et al., 2025), alignment with human values (Ouyang et al., 2022; Bai et al., 2022; Dai et al., 2023), and adherence to safety protocols (Mu et al., 2024; Huang et al., 2024; Zhang et al., 2024; Duan et al., 2024). Since these desired behaviors often differ significantly from those of the pretrained model (Ouyang et al., 2022; Bai et al., 2022; OpenAI, 2024), it is often assumed that achieving them requires substantial changes to the model’s parameters and therefore full finetuning is widely applied during RL (Cui et al., 2025; HuggingFace, 2025; Liu et al., 2025; Pan et al., 2025; Zeng et al., 2025). While RL with full finetuning is allowed to update all parameters, does it actually do so? This paper presents surprising findings and answers this question in the negative.
Finding 1 RL-induced parameter update sparsity in LLMs: RL updates only a small subnetwork of a pretrained large language model, leaving the rest of the parameters effectively unchanged.
Finding 1 is observed in all 7 widely-used RL algorithms studied, namely PPO (Schulman et al., 2017b), GRPO (Shao et al., 2024), ORPO (Hong et al., 2024), KTO (Ethayarajh et al., 2024), DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024) and PRIME (Cui et al., 2025), as well as supervised finetuning with rejection sampling (Xiong et al., 2025), and 10 models in our experiments, with the subnetworks consisting of as little as 5% of the model parameters in some cases (§3). It emerges intrinsically from RL finetuning, without any explicit sparsity-promoting regularization, architectural constraint, or use of parameter-efficient training or pruning methods. Moreover, we observe a strong consistency among the subnetworks emerged under different random seeds, training data and its order, and even different RL algorithms, suggesting that the pretrained model contains a partially transferable structure that is consistent across varied training conditions (§5).
Figure 1: Comparison in accumulated gradients in the SFT stage vs RL stage for popular released checkpoints. SFT stage has accumulated much denser updates, while RL is mostly sparse.
Interestingly, our experiments with PRIME suggest that approximately 20% of the parameters are consistently updated and make up the subnetwork. An additional 8% receive non-zero gradients during training that cancel out, while the remaining $\sim$70parameters remain untouched throughout the entire training process. This observation motivates us to articulate the following conjecture:
Figure 2: In PRIME, 72% parameters are never updated, 8% have gradients canceling each other out, and 20% constitute the subnetwork that is consistently updated (§6)
Conjecture 1 Fine-tuning only the subnetwork identified at the end of RL training, with all other parameters frozen, produces a model that is nearly identical to the original model that has undergone full finetuning, both in test accuracy and in parameter values.
More formally, let $\theta_{\text{full}}$ denote the parameters after full RL finetuning from the initial model $\theta_{\text{init}}$. Define a binary mask $m\in\{0,1\}^{|\theta_{\text{init}}|}$ where $m_{i}=1$ if $(\theta_{\text{init}}-\theta_{\text{full}})_{i}\neq 0$ and 00 otherwise. We finetune a second model from $\theta_{\text{init}}$ on the same data with the same hyperparameters and number of gradient updates, but, at each step, mask the gradients as $m\odot\nabla_{\theta}\mathcal{L}(\theta)$ right before the parameter update, so that only the identified subnetwork receives non-zero gradients. Let $\theta_{\text{sub}}$ denote the resulting parameters of this subnetwork finetuning process. Conjecture 1 states that $\theta_{\text{sub}}\approx\theta_{\text{full}}$. We provide supporting evidence for the conjecture on PRIME and DPO in §4.
We find that the updates of RL finetuning do not concentrate in specific layers or components of the transformer. Instead, nearly all parameter matrices receive similarly sparse updates (§3). An exception is the layer normalization layers, which receives little to no updates. Moreover, despite the sparsity in the updates, they are almost always full-rank. This suggests that, instead of forcing the updates to reside in a low-rank subspace as in LoRA (Hu et al., 2022), RL full finetuning updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent.
To better understand the potential reasons for this phenomenon, we conduct a series of experiments in §6. The results indicate that a primary factor is training on data that is near the policy distribution through, e.g., on-policy RL or performing supervised finetuning (SFT) on the same data before RL (Wang et al., 2024a; Cui et al., 2025). Intuitively, it requires less change to the policy distribution when the model learns on a sequence sampled from a distribution close to itself. In contrast, SFT often involves distribution shifts (Zhang et al., 2025) and densely updates the models in our experiments (§3). Other factors like KL-divergence regularization towards the reference model, gradient clipping (as used in PPO, GRPO, and PRIME), online vs. offline RL, all have limited impact on the sparsity of accumulated updates.
Our findings have profound implications on RL finetuning of LLMs. They suggest that RL concentrates optimization entirely on a small and consistently active subnetwork, while the remaining parameters remain effectively inert. Conjecture 1 goes beyond the lottery ticket hypothesis (LTH) (Frankle and Carbin, 2019): not only can the subnetwork, finetuned in isolation, match the performance of the full model in performance as posited by LTH, we show that it also converges to an effectively identical model. These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities compared to SFT (Chu et al., 2025; Setlur et al., 2025), potentially by updating substantially fewer parameters. They also open up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (Chen et al., 2022).
2 Related work and Background
2.1 Related Work
The Lottery Ticket Hypothesis (LTH; Frankle and Carbin, 2019) posited that dense neural networks contain sparse subnetworks capable of matching the performance of the full model when trained in isolation. Subsequent extensions to LLMs identified task-specific subnetworks that mitigate catastrophic forgetting without retraining entire models (Panda et al., 2024; Panigrahi et al., 2023; Yadav et al., 2023). Related efforts further discovered sparse subnetworks in pretrained language models crucial for encoding specific knowledge (Marks et al., 2025; Bayazit et al., 2024; Liu et al., 2022). Recent works have also explored exploiting the winning lotteries to improve training efficiency (Chen et al., 2022). While our observation is closely related to LTH, it differs in three core dimensions: (1) LTH identifies winning tickets by pruning, while we study subnetwork that naturally emerge; (2) LTH showed that the final model’s performance can be reproduced, we show that, in addition to the performance, the exact same model can almost be recovered; (3) LTH focuses on models trained from scratch, while we focus on finetuning from pretrained LLMs.
Sparse training methods exhibit notable benefits in RL efficiency (Graesser et al., 2022; Sidahmed et al., 2024; Sokar et al., 2022; Tan et al., 2023; Davelouis et al., 2025). Recent studies also employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) in RL and have achieved competitive performance alongside significantly reduced computational overhead (Sidahmed et al., 2024; Wang et al., 2025b). In contrast to approaches like LoRA that explicitly constrain updates to a small number of parameters, we find that fine-tuning the naturally emerging subnetwork can match or even surpass the performance of full-model finetuning. Moreover, despite its sparsity, the updates are nearly full-rank.
2.2 Background
We briefly introduce key concepts and notations to be used in onward discussion.
The sparsity of parameter updates. Let $\theta^{0},\theta^{1}\in\mathbb{R}^{n}$ denote the model parameters before and after finetuning, respectively. We define the update sparsity as $\text{sparsity}(\theta^{0},\theta^{1}):=1-\left\lVert\theta^{1}-\theta^{0}\right\rVert_{0}/n$, where $\left\lVert\cdot\right\rVert_{0}$ counts the number of non-zero elements. It is important to clarify that even the update $\theta_{1}-\theta_{0}$ is sparse, it does not imply that the finetuned model $\theta_{1}$ is sparse. Since no sparsity is assumed for $\theta_{0}$, a sparse update can still result in a dense $\theta_{1}$ if $\theta_{0}$ is dense.
Unless otherwise specified, we follow standard practice and consider two bfloat16 values as equal when their absolute difference does not exceed $10^{-5}$, to account for numerical precision limits. [^1] E.g., PyTorch uses $10^{-5}$ as the default tolerance for gradient checking. All models in our experiments are in the bfloat16 data type. Sparsity with different tolerance values can be found in Table 6 in the Appendices.
Learning from in-distribution data. We use “in-distribution” to refer to training on data drawn from a distribution that closely matches the current policy. An example is on-policy RL with, e.g., PPO, GRPO, and PRIME, which sample data online from the evolving policy during training. Another way to achieve in-distribution RL is to perform SFT on the same data used for subsequent RL, so that the policy adapts to the data distribution before RL. This is a common practice in off-policy methods like DPO and KTO. On-policy methods inherently train on in-distribution data, and off-policy methods can also do so when the training data closely matches the policy distribution. As we will show later in §6, training on in-distribution data is a primary reason for the update sparsity in RL.
KL-divergence regularization and gradient clipping in RL. Two widely adopted techniques to keep the policy from deviating too far from the reference model are KL-divergence regularization and gradient clipping. KL regularization (Schulman et al., 2017a), formally computed as $D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\mathbb{E}_{\pi_{\theta}}\left[\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right]$, constrains policy shifts. Gradient clipping further stabilizes training by bounding the update norm. Both are widely used in algorithms such as PPO, GRPO, and PRIME. In §6, we show that, counterintuitively, both have limited impact on the update sparsity.
3 RL Induces Sparse but Full-rank Updates; SFT Induces Dense Ones
This section aims to answer the following research question
  • RQ1:
    To what extent does RL induce sparse parameter updates and where in the model do these updates occur? How does SFT compare?
Setup. To answer this question, we analyze publicly released model checkpoints on Hugging Face released by the authors. With the exception of models where RL is applied directly to the pretrained base model (e.g., DeepSeek-R1-Zero), most models follow a conventional three-stage pipeline: pretraining, supervised fine-tuning (SFT), and RL. We analyze both the RL and SFT stages by measuring the update sparsity between model checkpoints before and after RL or SFT fine-tuning. Our experiments cover Tulu 8B/70(Lambert et al., 2025), Eurus 7B (Yuan et al., 2025; Cui et al., 2025), DeepSeek Math 7B (Shao et al., 2024), and KTO/SimPO models (Meng et al., 2024).
Table 1: Parameter update sparsity across different RL algorithms. We report sparsity for a suite of open models from Hugging Face. For all models, at least 68.5%—and often much more—of the parameters remain unchanged after RL.
Algo. Init Model RL Model Update Sparsity On-Policy KL Online DPO Llama-3.1-Tulu-3-8B-SFT Llama-3.1-Tulu-3-8B-DPO 81.4 Llama-3.1-Tulu-3-70B-SFT Llama-3.1-Tulu-3-70B-DPO 95.2 GRPO deepseek-math-7b-instruct deepseek-math-7b-rl 68.5 DeepSeek v3 base DeepSeek-R1-Zero 86.0 ORPO mistral-7B-v0.1 mistral-orpo-beta 76.9 KTO Eurus-7b-sft Eurus-7b-kto 96.0 Llama-3-Base-8B-SFT Llama-3-Base-8B-SFT-KTO 81.2 PPO mistral-7b-sft math-shepherd-mistral-7b-rl 80.8 SimPO Meta-Llama-3-8B-Instruct Llama-3-Instruct-8B-SimPO 86.5 PRIME Eurus-2-7b-sft Eurus-2-7B-PRIME 77.0
Results. As shown in Table 1, for all RL-finetuned models, 68.5%–96.0% of parameters remain unchanged after RL. This trend holds across different RL algorithms and model families. Particularly, Deepseek-R1-Zero presents a update sparsity of 86.0%, regardless of directly training from the pretrained base model, namely RL-Zero (DeepSeek-AI et al., 2025), and large-scale training for over 8K steps. Although exact training configurations are not always available, we observe that within the same model family, larger models tend to show higher sparsity. Importantly, all of these models are trained using full finetuning without any sparsity-promoting regularization techniques or constraints. This suggests that the update sparsity is intrinsic to RL training rather than a result of design artifacts.
In contrast, Figure 1 shows that SFT induces dense updates (only 6%-15% sparsity). These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities than SFT (Chu et al., 2025; Setlur et al., 2025), possibly by updating substantially fewer parameters.
Takeaway 1 RL leads to consistently sparse parameter updates (often $>$70% sparsity) while SFT produces dense updates. Sparsity emerges as a characteristic property of RL finetuning in LLMs.
Figure 3: Layerwise and per-parameter-matrix update sparsity for DPO (left) and PRIME (right). All layers are similarly sparsely updated, with the only exception of the layer normalization layers, which receive little to no updates.
Table 2: Mean ranks of update matrices, as a percentage of maximum possible rank across models after RL finetuning.
Model and Algo. Update Rank (%)
Tulu 8B (DPO) 99.8
Eurus 7B (PRIME) 99.5
Llama-3 8B (KTO) 99.2
DeepSeek Math 7B (GRPO) 99.4
Almost all transformer layers receive similarly sparse updates. We next examine how parameter updates in RL are distributed across the model layers and individual parameter matrices (e.g., Q, K, V projections), based on DPO and PRIME models. If updates were concentrated in a subset of layers or modules, one could exploit that structure for enhancing the efficiency (Pan et al., 2024). Figure 3 shows layerwise and per-matrix sparsity across the models. The "Average Sparsity" is over each transformer layer, while others correspond to specific parameter matrices. We observe that parameter updates are distributed across different matrices rather than localized to specific ones. Except for consistently high sparsity in layer normalization layers, most layers exhibit similar sparsity levels. Our results show that sparsity is relatively even across the model. This suggests that recovering the behavior of the fully finetuned model requires updating all layers, albeit with only a subset of parameters in each.
Updates are sparse but full-rank. Given the sparsity of RL-induced updates, a natural question is whether these updates are also low-rank. This distinction between low-rank and sparse updates is important: the former would imply that finetuning operates within a subspace, while the latter implies that a small subset of parameters (that can span the full parameter space) are selected to finetune. Notably, while the updates are sparse, a closer inspection reveals that they are nearly full rank (Tab 2). To compute rank, we calculate the average rank of individual update matrices across all layers. We further examine the rank of the update for each layer and parameter matrix, and find that most are full-rank throughout the model. These findings suggest that RL updates are localized to a subset of the parameters that almost span the full subspaces that the parameter matrices can represent, instead of residing in a low-rank subspace.
Takeaway 2 All layers and parameter matrices receive similarly sparse but full-rank updates. While layer normalization parameters are almost never updated.
4 Finetuning the Subnetwork Alone Can Reproduce the Full-finetuned Model
Since RL primarily fine-tunes a small subnetwork, we investigate two research questions inspired by but extending beyond the Lottery Ticket Hypothesis (LTH):
  • RQ2:
    Can finetuning the subnetwork in isolation recover the performance of the full-finetuned model?
  • RQ3:
    Can subnetwork-only finetuning also recover the exact parameter values produced by full RL finetuning? This section answers both in the positive.
Setup. We follow the procedure described in §1 to obtain two models: one with full finetuning $\theta_{\text{full}}$, and another finetuned on the same data and hyperparameters but updating only the subnetwork $\theta_{\text{sub}}$. We experiment on two very different algorithms DPO, an off-policy algorithm using implicit outcome rewards, and PRIME, an on-policy one with process reward models, to ensure that our conclusion can generalize. We implement DPO with Open-Instruct and PRIME with verl. The exact hyperparameter choices for both can be found in Appendix B. For evaluation, we choose a subset of tasks reported in the original papers for both. For DPO we choose the LSAT (Wang et al., 2022), LogiQA (Liu et al., 2021) splits from AGIEval (Zhong et al., 2024), Math split of MMLU Pro (Wang et al., 2024b). For PRIME, we report results on the MATH500 (Hendrycks et al., 2021) benchmark across difficulty levels. For evaluation in DPO we use olmes. [^2] Open-Instruct: https://github.com/allenai/open-instruct; verl: https://github.com/volcengine/verl/; olmes: https://github.com/allenai/olmes/
Figure 4: Training loss for training DPO with subnetwork finetuning and full finetuning. Training the subnetwork in isolation consistently causes train loss to be lower.
Results. In DPO, 94.0% weights are same between $\theta_{\text{full}}$ and $\theta_{\text{sub}}$; it is 90.5% for PRIME. Notably, for both DPO and PRIME, $\theta_{\text{full}}$ and $\theta_{\text{sub}}$ are 100% identical when using a tolerance of $10^{-4}$ instead of the default $10^{-5}$, indicating that the two models converge to nearly identical parameter values. As shown in Tables 3(a) and 3(b), $\theta_{\text{sub}}$ matches or outperforms $\theta_{\text{full}}$ on all tasks across both algorithms. These results suggest that the parameters outside the subnetwork play little role in the optimization process, and freezing them has a negligible or even beneficial impact on the model’s performance. We further observe that the training loss is consistently lower in the subnetwork finetuning setting than full finetuning (Fig. 4). They also provide supporting evidence for our Conjecture 1 in §1. This finding opens up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (Chen et al., 2022).
Table 3: Test set performance of $\theta_{\text{full}}$ and $\theta_{\text{sub}}$ trained with DPO and PRIME. Training only the subnetwork ($\theta_{\text{sub}}$) can achieve better performance than full finetuning ($\theta_{\text{full}}$). Lvl. indicates the difficulty levels of MATH500.
6cm][t]0.48 Task $\theta_{\text{full}}$ $\theta_{\text{sub}}$ $\Delta$ AGIEval LSAT-AR 21.3 24.8 +3.5 AGIEval LSAT-LR 53.1 54.7 +1.6 AIGEval LogiQA-en 43.5 45.5 +2.0 GPQA 32.8 32.8 +0.0 MMLU Pro Math 50.8 51.6 +0.8 Avg 40.3 41.9 +1.6
(a) DPO
6cm][t]0.48 Lvl. $\theta_{\text{full}}$ $\theta_{\text{sub}}$ $\Delta$ 1 93.0 93.0 +0.0 2 85.6 85.6 +0.0 3 82.9 83.8 +0.9 4 71.1 74.2 +3.1 5 40.3 45.5 +5.2 Overall 69.8 72.2 +2.4
(b) PRIME
5 Consistency of Subnetworks Across Seeds, Data, and Algorithms
This section aims to answer the following research question:
  • RQ4:
    How consistent is the RL-updated subnetwork under varying training conditions such as random seed, training data, RL algorithm, and even all of them?
If the subnetwork remains largely consistent across these variations, it would suggest that the identified subnetwork is not merely an artifact of specific training configuration but a generalizable and transferable structure of the pretrained model.
Setup. To quantify the similarity between two subnetworks, we define an overlap metric. Let $s_{1}$ and $s_{2}$ denote the sparsity levels of two models, and let $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ be the sets of indices of the updated parameters. The size of the common subnetwork is given by $|\mathcal{I}_{1}\cap\mathcal{I}_{2}|$. One-sided overlap is then $o_{1}=|\mathcal{I}_{1}\cap\mathcal{I}_{2}|/|\mathcal{I}_{1}|=|\mathcal{I}_{1}\cap\mathcal{I}_{2}|/(1-s_{1})$, which quantifies the proportion of $\mathcal{I}_{1}$ that is covered by the subnetwork $\mathcal{I}_{2}$, i.e., how well $\mathcal{I}_{2}$ captures the parameters updated in $\mathcal{I}_{1}$. Similarly, $o_{2}=|\mathcal{I}_{1}\cap\mathcal{I}_{2}|/(1-s_{2})$ quantifies how well $\mathcal{I}_{1}$ captures the parameters updated in $\mathcal{I}_{2}$. We compare the observed overlaps $o_{1}$ and $o_{2}$ against a random guessing baseline, where a subnetwork is constructed by uniformly selecting the same number of parameters as identified by RL (Appendix E).
We evaluate three settings: (1) varying the random seed alone, (2) varying the training data alone, and (3) changing the seed, data, and the RL algorithm altogether as a stress test. We conduct controlled experiments and all factors not under investigation are the same. Unless otherwise mentioned, all ablations were done with a batch size of 32, trained for one epoch with the base model Tulu-3-8B-SFT. When varying the training data, we switch between the Tulu preference dataset333allenai/llama-3.1-tulu-3-8b-preference-mixture and the PRIME rollout dataset444 PRIME-RL/Eurus-2-Rollout. It has model generations for math datasets alongside a label for correctness. To adapt the rollout dataset to DPO format, we select only positive samples, and pair it with a randomly sampled negative one. When varying the RL algorithm, we train a DPO model initialized from PRIME-RL/Eurus-2-7B-SFT and compare it to the PRIME-RL/Eurus-2-7B-PRIME model.
Results. Table 4 reports our observed overlap. Despite changes in initialization, the resulting subnetworks show substantial overlap—well above the random baseline. For instance, varying the random seed yields overlaps of $o_{1}=60.5\%$ and $o_{2}=60.6\%$. Similar consistency is observed when the training dataset is varied. Even under a stress test, where the data, seed, and RL algorithm are all changed, we still observe notable overlaps of 59.1% and 33.2%. These findings indicate the presence of a subnetwork that is at least partially transferrable to other different settings.
Table 4: Subnetwork overlap varying random seeds, training data, and RL algorithms. Despite these changes, subnetworks show non-trivial overlap compared to random-guessing baselines.
Variation Axis Setting Random RL Subnetwork Sparsity
Seed $\mathcal{I}_{1}$: 42 $o_{1}$: 36.7 $o_{1}$: 60.5 63.3
$\mathcal{I}_{2}$: 123 $o_{2}$: 36.7 $o_{2}$: 60.6 63.3
Data $\mathcal{I}_{1}$: Tulu Data $o_{1}$: 14.6 $o_{1}$: 26.7 63.3
$\mathcal{I}_{2}$: PRIME Data $o_{2}$: 36.7 $o_{2}$: 67.1 85.4
Seed + Data + Algo. $\mathcal{I}_{1}$: DPO $o_{1}$: 23.0 $o_{1}$: 59.1 87.1
$\mathcal{I}_{2}$: PRIME $o_{2}$: 12.9 $o_{2}$: 33.2 77.0
Takeaway 3 For a given base model, we observe substantially higher subnetwork overlap than random guessing across different seeds, training data, and RL algorithms. This suggests the potential of a consistent and at least partially transferrable subnetwork structure across these different training settings.
While the observed subnetwork overlap across seeds, datasets, and training algorithms falls short of 100%, it suggests that partial subnetwork reuse may still offer practical utility. In particular, partial subnetwork reuse could reduce redundant computation across repeated RL runs, such as those in hyperparameter sweeps or ablation studies, by partially reusing the subnetworks. In addition, one might be able to reuse part of the subnetwork identified by a cheaper algorithm like DPO and reuse it in more expensive ones like PPO, greatly reducing the training cost without sacrificing performance.
6 Why Do the Subnetworks Emerge?
This section answers investigates the following research question:
  • RQ5:
    What factors contribute to the update sparsity observed in RL finetuning?
We investigate the following factors: gradient clipping, KL regularization towards a reference policy, performing SFT prior to RL, and the number of RL update steps. Our investigation suggests that the dominating factor is how close the training data distribution is to the policy’s, i.e., whether the training data is in-distribution. Another important factor is the total number of gradient updates.
Figure 5: Update sparsity of intermediate checkpoints of a training run of PRIME. We observe that with more training the sparsity slowly decays.
Gradient clipping and KL regularization. As discussed in §2.2, gradient clipping and KL regularization are commonly used to keep the policy from deviating too far away from a reference model. Since both mechanisms explicitly suppress large parameter updates, they are natural candidates for contributing factors to the observed update sparsity. To test their impact, we train a GRPO variant using Qwen-2.5-7B-Instruct, comparing models with and without these regularization terms. We find that both configurations exhibit comparable sparsity levels, suggesting that neither gradient clipping nor KL regularization is a primary driver of the update sparsity. In our experiments, the GRPO variant trained with KL regularization achieved a sparsity of 69.8%, while the variant trained without KL regularization reached 68.8%. Further, SimPO removes the KL term by dropping the base policy normalization in DPO, and as reported in 1 SimPO also produces sparse updates, providing further negative evidence for KL.
Performing SFT before RL. A common design choice is to perform SFT on the same data as the subsequent RL Ouyang et al. (2022). However, as shown in Table 1, our findings extend to models such as DeepSeek-R1-Zero, which forgoes SFT entirely yet still exhibit high update sparsity. This suggests that SFT is not a main contributing factor to the update sparsity in RL finetuning.
Training duration. It is intuitive that with more gradient steps, a model is expected to drift more from the base model. Figure 5 shows how update sparsity changes during the training of PRIME. As training progresses, update sparsity gradually decreases but eventually shows signs of convergence to around 80%. Notably, DeepSeek-R1-Zero undergoes 8K training steps (numbers from Figure 2 in DeepSeek-AI et al. (2025)) using GRPO, over 20$\times$ more than PRIME, but shows a comparable update sparsity (86%). Therefore, we conjecture that training duration’s impact on update sparsity is more prominent during early training but gradually decreases as training progresses.
Figure 6: Percentage of updated weights that are outside the final subnetwork across training steps
Figure 6 shows the percentage of updated parameters (relative to model size) that lie outside the final subnetwork. This proportion increases during the early stages of training but steadily declines in later stages. This trend suggests that some parameters outside the final subnetwork receive non-zero gradient updates that cancel out. Overall, about 8.5% of parameters that are ever updated during training fall outside the final subnetwork.
While it is possible, though less likely, that all models we study, including DeepSeek-R1-Zero (86.0% update sparsity after 8K steps), are severely undertrained, and that the observed update sparsity would diminish with substantially more training. Nonetheless, we question the practicality of this hypothesis since it runs counter to the RL literature arguing against overtraining to prevent overfitting and improve generalization Fu et al. (2019).
Training on in-distribution data. Intuitively, when gradients are computed on sequences that the policy already assigns high probabilities to, little update to the parameters would be needed. We evaluated two scenarios: (1) rejection sampling, and (2) DPO on out-of-distribution data by not performing SFT pridor to RL. Since we’ve already learned from §3 that DPO with in-distribution data induces sparse updates while SFT (with out-of-distribution data) induces dense ones, this additional experiment can serve as a control group to isolate the factor of training on in- vs. out-of-distribution data.
As shown in Table 5, Our experiments reveal that SFT on in-distribution data produces sparse updates, while DPO with out-of-distribution data produces dense ones. Specifically, performing SFT with Qwen/Qwen2.5-Math-7B on rejection sampled in-distribution data yields around 90.0% update sparsity. This is reinforced by the examination of a previous work: RAFT++ (Xiong et al., 2024), which performs supervised finetuning with iterative rejection sampling, yields an update sparsity of 69.4%. In contrast, DPO on out-of-distribution data produces dense updates in zephyr-7b-beta models, with a 6.8% update sparsity. These findings suggest that training on in-distribution data is a major driver of update sparsity in not only RL, but also SFT.
Takeaway 4 Training on in-distribution data is a primary reason of update sparsity; KL-divergence regularization and gradient clipping have limited impact.
Table 5: RFT indicates rejection-sampling fine-tuning (Touvron et al., 2023; Dong et al., 2023), and RAFT++ is iterative RFT A comparative analysis across SFT and DPO as well as in- vs out-of-distribution training shows that in-distribution consistently produces sparse updates.
Model Method Sparsity (%) SFT/RL In-Dist
Qwen2.5-Math-7B RFT 91.2 SFT
Qwen2.5-Math-7B RAFT++ 69.4 SFT
Llama-3.1-8B-SFT SFT 6.8 SFT
Llama-3.1-8B-SFT DPO 6.8 RL
Zephyr-7b-Beta DPO 7.7 RL
Llama-3.1-8B-DPO DPO 81.4 RL
7 Limitations and Future Work
Because RL is computationally demanding, we choose to vary one factor at a time; yet the observed sparsity may actually result from complex interactions among many, an avenue future work should examine. Further, fully controlled experiments are computationally prohibitive, and we thus sometimes resort to resort to public checkpoints. While our experiments focus on language models, it would be interesting to explore the same questions for multimodal and diffusion models. Subsequent research could investigate methods for early identification of the sparse subnetwork and ways to leverage its structure for more efficient learning. Finally, our empirical findings invite a deeper theoretical analysis, with the goal of uncovering theoretical explanations for the update sparsity in RL.
8 Conclusion
Our study reveals that RL finetuning in LLMs updates only a sparse subnetwork constituting approximately 5%-30% of total parameters, leaving the rest unchanged. This sparsity emerges intrinsically, without explicit spasity promoting techniques such as regularization or structural constraints. Crucially, finetuning just this subnetwork in isolation reproduces the full model’s performance, aligning closely with original parameter values. For a given base model across different seeds, datasets and learning algorithms, a non-trivial portion of the subnetwork remains the same. Our findings highlight that learning from in-distribution samples while training is a key driver of this phenomenon, pointing towards more efficient and effective training strategies in RL-based finetuning of LLMs.
Appendix A Appendix
Algorithm Model (RL Checkpoint) Tol = 1e-8 Tol = 1e-7 Tol = 1e-6 Tol = 1e-5
DPO allenai/Llama-3.1-Tulu-3-8B-DPO 76.04 76.04 76.14 81.38
DPO allenai/Llama-3.1-Tulu-3-70B-DPO 87.58 87.59 87.79 95.24
GRPO deepseek-ai/deepseek-math-7b-rl 68.14 68.14 68.14 68.53
ORPO kaist-ai/mistral-orpo-beta 73.16 73.18 73.23 76.94
ORPO kaist-ai/mistral-orpo-alpha 50.40 50.41 50.48 53.23
KTO openbmb/Eurus-7b-kto 71.78 71.79 73.14 95.98
PPO peiyi9979/math-shepherd-mistral-7b-rl 52.45 52.47 53.21 80.77
PPO PRIME-RL/Eurus-2-7B-PRIME 75.26 75.27 75.36 77.04
SimPO Llama-3-Instruct-8B-SimPO 71.00 71.00 71.10 76.42
SimPO Llama-3-Base-8B-SFT-SimPO 79.47 79.47 79.60 86.52
SimPO Mistral-7B-Instruct-SimPO 59.37 59.40 60.31 89.07
SimPO Mistral-7B-Base-SFT-SimPO 62.58 62.60 63.56 91.44
Table 6: Sparsity (%) of parameter updates under different thresholds across RL algorithms and RL checkpoints.
Appendix B Hyperparameter choices for Gradient Masking experiments
DPO: For DPO, we fine-tuned the LLaMA-3.1-Tulu-3-8B model using Direct Preference Optimization (DPO) with bfloat16 mixed-precision and DeepSpeed Stage 3 for memory and compute efficiency across 8 processes. Training uses a sequence length of 2048 tokens with an effective batch size of 128, achieved by setting the per-device batch size to 1 with 16 gradient accumulation steps. A linear learning rate schedule is applied with a peak learning rate of $5\times 10^{-7}$ and a warmup ratio of 0.1, without weight decay. The model is trained for one epoch on the allenai/llama-3.1-tulu-3-8b-preference-mixture dataset.
PRIME: For PRIME, We fine-tune Qwen2.5-Math-7B using on a mixture of GSM8K and MATH datasets. The training batch size is set to 64. The actor is optimized with a learning rate of $5\times 10^{-7}$ while the reward model is trained with a learning rate of $1\times 10^{-6}$. We performed four rollouts are performed per sample. We use gradient clipping of 10.0, and a temperature $\beta$ of 0.05. Training is conducted on for 15 epochs.
Appendix C Model Checkpoints: SFT vs RL sparsity comparison
SFT Checkpoints.
We compare the following base and SFT checkpoints:
  • meta-llama/Llama-3.1-8B vs. allenai/Llama-3.1-Tulu-3-8B-SFT
  • meta-llama/Llama-3.1-70B vs. allenai/Llama-3.1-Tulu-3-70B-SFT
  • Qwen/Qwen2.5-Math-7B vs. PRIME-RL/Eurus-2-7B-SFT
RL Checkpoints.
We compare the following SFT and RL-finetuned checkpoints:
  • allenai/Llama-3.1-Tulu-3-8B-SFT vs. allenai/Llama-3.1-Tulu-3-8B-DPO
  • allenai/Llama-3.1-Tulu-3-70B-SFT vs. allenai/Llama-3.1-Tulu-3-70B-DPO
  • PRIME-RL/Eurus-2-7B-SFT vs. PRIME-RL/Eurus-2-7B-PRIME
Appendix D Training Dynamics
We analyzed intermediate checkpoints of the Prime model. Notably, our goal is to observe the convergence of the sparsity with training time. Does the sparsity decay with gradient steps ? Does it asymptotically reach a sparsity level of zero or is the convergence to a non-zero point ?
Experimental setup
We analyze 21 intermediate checkpoints from a training run of the PRIME model. Let the model parameters at these checkpoints be denoted by $\theta_{1},\theta_{2},\ldots,\theta_{21}$, and let $\theta_{\text{init}}$ denote the parameters of the corresponding base model (i.e., PRIME-RL/Eurus-7b-sft). We define the sparsity between the base model to checkpoint $k$ as $sparsity_{k}$ = sparsity($\theta_{k},\theta_{init}$), and the sparsity between two checkpoints $i$ and $j$ as $sparsity_{ij}$ = sparsity($\theta_{j},\theta_{i}$).
Key Findings
Our analysis begins by examining how the $sparsity_{k}$ evolve with training progress, offering insight into the update patterns.
Figure 7: Sparsity Analysis of intermediate checkpoints of Prime (a) shows the sparsity
Figure 7(a) illustrates the sparsity of intermediate checkpoints alongside their accuracy on the MATH500 and AIME2024 tasks. The plot clearly demonstrates that all intermediate checkpoints exhibit non-trivial sparsity. Furthermore, as training progresses, the sparsity converges to a numerically significant asymptote, suggesting that a substantial proportion of weights remain unaffected even after prolonged period of training. In Figure 7(b), we report the sparsity of the $sparsity_{ij}$ for all consecutive checkpoint pairs, i.e., where $j=i+1$. Notably, in each successive step, on average, only 7% of the weights receive a non-zero gradient update.
Appendix E Random Guessing baseline
If model 1 has sparsity $s_{1}$ and model 2 has sparsity $s_{2}$, the expected overlap is given by: $\frac{(1-s_{1})\cdot(1-s_{2})}{100}$. Normalizing like earlier, we get $\mathcal{O}_{1,random}=\frac{(1-s_{2})}{100}$ and $\mathcal{O}_{2,random}=\frac{(1-s_{1})}{100}$. i.e. for random guessing the overlap for model 1 is the density of model 2.

### v2

@@
 It is observed across all 7 widely-used RL algorithms (e.g., PPO, GRPO, DPO) and all 10 LLMs from different families in our experiments.
-This sparsity is intrinsic and occurs without any explicit sparsity-promoting regularizations or architectural constraints.
+This sparsity occurs without any explicit sparsity-promoting regularizations or architectural constraints.
 Finetuning the subnetwork alone recovers the test accuracy, and, remarkably, produces a model nearly identical to the one obtained via full finetuning.
@@
 We conjecture that the this update sparsity can be primarily attributed to training on data that is near the policy distribution; techniques that encourage the policy to remain close to the pretrained model, such as the KL regularization and gradient clipping, have limited impact.
+Our code is available at https://github.com/SagnikMukherjee/sparsity_in_rl.
 1 Introduction
-Reinforcement learning (RL) (Sutton et al., 1998; Ouyang et al., 2022; Ziegler et al., 2020; Ramamurthy et al., 2023; Sun et al., 2024; Zhou et al., 2025) is an important post-pretraining stage for adapting large language models (LLMs) to solving complex reasoning problems (Lightman et al., 2023; Wang et al., 2025a, c; Cui et al., 2025), alignment with human values (Ouyang et al., 2022; Bai et al., 2022; Dai et al., 2023), and adherence to safety protocols (Mu et al., 2024; Huang et al., 2024; Zhang et al., 2024; Duan et al., 2024).
-Since these desired behaviors often differ significantly from those of the pretrained model (Ouyang et al., 2022; Bai et al., 2022; OpenAI, 2024), it is often assumed that achieving them requires substantial changes to the model’s parameters and therefore full finetuning is widely applied during RL (Cui et al., 2025; HuggingFace, 2025; Liu et al., 2025; Pan et al., 2025; Zeng et al., 2025).
+Reinforcement learning (RL) (sutton1998reinforcement; ouyang2022traininglanguagemodelsfollow; ziegler2020finetuninglanguagemodelshuman; ramamurthy2023is; sun-etal-2024-aligning; zhou2025reinforcedmllmsurveyrlbased) is an important post-pretraining stage for adapting large language models (LLMs) to solving complex reasoning problems (lightman2023letsverifystepstep; wang2025otcoptimaltoolcalls; wang2025reinforcementlearningreasoninglarge; cui2025processreinforcementimplicitrewards), alignment with human values (ouyang2022traininglanguagemodelsfollow; bai2022constitutionalaiharmlessnessai; dai2023saferlhfsafereinforcement), and adherence to safety protocols (mu2024rulebasedrewardslanguage; huang2024oneshot; zhang2024negative; DBLP:journals/corr/abs-2403-03419).
+Since these desired behaviors often differ significantly from those of the pretrained model (ouyang2022traininglanguagemodelsfollow; bai2022constitutionalaiharmlessnessai; OpenAIOS), it is often assumed that achieving them requires substantial changes to the model’s parameters and therefore full finetuning is widely applied during RL (cui2025processreinforcementimplicitrewards; openr1; liu2025understanding; tinyzero; zeng2025simplerlzooinvestigatingtamingzero).
+Finding 1 RL-induced parameter update sparsity in LLMs: RL updates only a small subnetwork of a pretrained large language model, leaving the rest of the parameters effectively unchanged.
 While RL with full finetuning is allowed to update all parameters, does it actually do so?
 This paper presents surprising findings and answers this question in the negative.
-Finding 1 RL-induced parameter update sparsity in LLMs: RL updates only a small subnetwork of a pretrained large language model, leaving the rest of the parameters effectively unchanged.
-Finding 1 is observed in all 7 widely-used RL algorithms studied, namely PPO (Schulman et al., 2017b), GRPO (Shao et al., 2024), ORPO (Hong et al., 2024), KTO (Ethayarajh et al., 2024), DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024) and PRIME (Cui et al., 2025), as well as supervised finetuning with rejection sampling (Xiong et al., 2025), and 10 models in our experiments, with the subnetworks consisting of as little as 5% of the model parameters in some cases (§3).
-It emerges intrinsically from RL finetuning, without any explicit sparsity-promoting regularization, architectural constraint, or use of parameter-efficient training or pruning methods.
+Finding 1 is observed in all 7 widely-used RL algorithms studied, namely PPO (schulman2017proximalpolicyoptimizationalgorithms), GRPO (shao2024deepseekmathpushinglimitsmathematical), ORPO (hong2024orpomonolithicpreferenceoptimization), KTO (ethayarajh2024ktomodelalignmentprospect), DPO (rafailov2023direct), SimPO (meng2024simposimplepreferenceoptimization) and PRIME (cui2025processreinforcementimplicitrewards), as well as supervised finetuning with rejection sampling (xiong2025minimalistapproachllmreasoning), and 10 models in our experiments, with the subnetworks consisting of as little as 5% of the model parameters in some cases (§3).
+It emerges without any explicit sparsity-promoting regularization, architectural constraint, or use of parameter-efficient training or pruning methods.
 Moreover, we observe a strong consistency among the subnetworks emerged under different random seeds, training data and its order, and even different RL algorithms, suggesting that the pretrained model contains a partially transferable structure that is consistent across varied training conditions (§5).
@@
 Interestingly, our experiments with PRIME suggest that approximately 20% of the parameters are consistently updated and make up the subnetwork.
-An additional 8% receive non-zero gradients during training that cancel out, while the remaining $\sim$70parameters remain untouched throughout the entire training process.
+An additional 8% receive non-zero gradients during training that cancel out, while the remaining $\sim$70% parameters remain untouched throughout the entire training process.
 This observation motivates us to articulate the following conjecture:
@@
 More formally, let $\theta_{\text{full}}$ denote the parameters after full RL finetuning from the initial model $\theta_{\text{init}}$.
-Define a binary mask $m\in\{0,1\}^{|\theta_{\text{init}}|}$ where $m_{i}=1$ if $(\theta_{\text{init}}-\theta_{\text{full}})_{i}\neq 0$ and 00 otherwise.
+Define a binary mask $m\in\{0,1\}^{|\theta_{\text{init}}|}$ where $m_{i}=1$ if $(\theta_{\text{init}}-\theta_{\text{full}})_{i}\neq 0$ and 0 otherwise.
 We finetune a second model from $\theta_{\text{init}}$ on the same data with the same hyperparameters and number of gradient updates, but, at each step, mask the gradients as $m\odot\nabla_{\theta}\mathcal{L}(\theta)$ right before the parameter update, so that only the identified subnetwork receives non-zero gradients.
@@
 Moreover, despite the sparsity in the updates, they are almost always full-rank.
-This suggests that, instead of forcing the updates to reside in a low-rank subspace as in LoRA (Hu et al., 2022), RL full finetuning updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent.
+This suggests that, instead of forcing the updates to reside in a low-rank subspace as in LoRA (hu2022lora), RL full finetuning updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent.
 To better understand the potential reasons for this phenomenon, we conduct a series of experiments in §6.
-The results indicate that a primary factor is training on data that is near the policy distribution through, e.g., on-policy RL or performing supervised finetuning (SFT) on the same data before RL (Wang et al., 2024a; Cui et al., 2025).
+The results indicate that a primary factor is training on data that is near the policy distribution through, e.g., on-policy RL or performing supervised finetuning (SFT) on the same data before RL (wang2024mathshepherdverifyreinforcellms; cui2025processreinforcementimplicitrewards).
 Intuitively, it requires less change to the policy distribution when the model learns on a sequence sampled from a distribution close to itself.
-In contrast, SFT often involves distribution shifts (Zhang et al., 2025) and densely updates the models in our experiments (§3).
+In contrast, SFT often involves distribution shifts (zhang2025bestinstructiontuningdatafit) and densely updates the models in our experiments (§3).
 Other factors like KL-divergence regularization towards the reference model, gradient clipping (as used in PPO, GRPO, and PRIME), online vs.
 offline RL, all have limited impact on the sparsity of accumulated updates.
-Our findings have profound implications on RL finetuning of LLMs.
-They suggest that RL concentrates optimization entirely on a small and consistently active subnetwork, while the remaining parameters remain effectively inert.
-Conjecture 1 goes beyond the lottery ticket hypothesis (LTH) (Frankle and Carbin, 2019): not only can the subnetwork, finetuned in isolation, match the performance of the full model in performance as posited by LTH, we show that it also converges to an effectively identical model.
-These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities compared to SFT (Chu et al., 2025; Setlur et al., 2025), potentially by updating substantially fewer parameters.
-They also open up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (Chen et al., 2022).
+Our findings have important implications for the RL fine-tuning stage of LLMs.
+They suggest that when RL fine-tuning is performed on data closely aligned with the current policy, as is typical in practice, optimization concentrates primarily on a small, consistently active subnetwork, leaving most other parameters effectively inert.
+Conjecture 1 goes beyond the lottery ticket hypothesis (LTH) (frankle2018the): not only can the subnetwork, finetuned in isolation, match the performance of the full model in performance as posited by LTH, we show that it also converges to an effectively identical model.
+These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities compared to SFT (chu2025sftmemorizesrlgeneralizes; setlur2025scalingtesttimecomputeverification), potentially by updating substantially fewer parameters.
+They also open up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (chen2022coarseninggranularitystructurallysparse).
 2 Related work and Background
 2.1 Related Work
-The Lottery Ticket Hypothesis (LTH; Frankle and Carbin, 2019) posited that dense neural networks contain sparse subnetworks capable of matching the performance of the full model when trained in isolation.
-Subsequent extensions to LLMs identified task-specific subnetworks that mitigate catastrophic forgetting without retraining entire models (Panda et al., 2024; Panigrahi et al., 2023; Yadav et al., 2023).
-Related efforts further discovered sparse subnetworks in pretrained language models crucial for encoding specific knowledge (Marks et al., 2025; Bayazit et al., 2024; Liu et al., 2022).
-Recent works have also explored exploiting the winning lotteries to improve training efficiency (Chen et al., 2022).
+The Lottery Ticket Hypothesis (LTH; frankle2018the) posited that dense neural networks contain sparse subnetworks capable of matching the performance of the full model when trained in isolation.
+Subsequent extensions to LLMs identified task-specific subnetworks that mitigate catastrophic forgetting without retraining entire models (panda2024lotteryticketadaptationmitigating; pmlr-v202-panigrahi23a; yadav2023tiesmergingresolvinginterferencemerging).
+Related efforts further discovered sparse subnetworks in pretrained language models crucial for encoding specific knowledge (marks2025sparsefeaturecircuitsdiscovering; bayazit2024discoveringknowledgecriticalsubnetworkspretrained; liu2022winwindealsparserobust).
+Recent works have also explored exploiting the winning lotteries to improve training efficiency (chen2022coarseninggranularitystructurallysparse).
 While our observation is closely related to LTH, it differs in three core dimensions: (1) LTH identifies winning tickets by pruning, while we study subnetwork that naturally emerge; (2) LTH showed that the final model’s performance can be reproduced, we show that, in addition to the performance, the exact same model can almost be recovered; (3) LTH focuses on models trained from scratch, while we focus on finetuning from pretrained LLMs.
-Sparse training methods exhibit notable benefits in RL efficiency (Graesser et al., 2022; Sidahmed et al., 2024; Sokar et al., 2022; Tan et al., 2023; Davelouis et al., 2025).
-Recent studies also employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) in RL and have achieved competitive performance alongside significantly reduced computational overhead (Sidahmed et al., 2024; Wang et al., 2025b).
+Sparse training methods exhibit notable benefits in RL efficiency (graesser2022statesparsetrainingdeep; sidahmed2024parameterefficientreinforcementlearning; sokar2022dynamicsparsetrainingdeep; tan2023rlx2trainingsparsedeep; davelouis2025interplaysparsitytrainingdeep).
+Recent studies also employ Low-Rank Adaptation (LoRA) (hu2022lora) in RL and have achieved competitive performance alongside significantly reduced computational overhead (sidahmed2024parameterefficientreinforcementlearning; wang2025tinatinyreasoningmodels).
 In contrast to approaches like LoRA that explicitly constrain updates to a small number of parameters, we find that fine-tuning the naturally emerging subnetwork can match or even surpass the performance of full-model finetuning.
@@
 Two widely adopted techniques to keep the policy from deviating too far from the reference model are KL-divergence regularization and gradient clipping.
-KL regularization (Schulman et al., 2017a), formally computed as $D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\mathbb{E}_{\pi_{\theta}}\left[\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right]$, constrains policy shifts.
+KL regularization (schulman2017trustregionpolicyoptimization), formally computed as $D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\mathbb{E}_{\pi_{\theta}}\left[\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right]$, constrains policy shifts.
 Gradient clipping further stabilizes training by bounding the update norm.
@@
 We analyze both the RL and SFT stages by measuring the update sparsity between model checkpoints before and after RL or SFT fine-tuning.
-Our experiments cover Tulu 8B/70(Lambert et al., 2025), Eurus 7B (Yuan et al., 2025; Cui et al., 2025), DeepSeek Math 7B (Shao et al., 2024), and KTO/SimPO models (Meng et al., 2024).
+Our experiments cover Tulu 8B/70(lambert2025tulu3pushingfrontiers), Eurus 7B (yuan2024advancing; cui2025processreinforcementimplicitrewards), DeepSeek Math 7B (shao2024deepseekmathpushinglimitsmathematical), and KTO/SimPO models (meng2024simposimplepreferenceoptimization).
 Table 1: Parameter update sparsity across different RL algorithms.
@@
 This trend holds across different RL algorithms and model families.
-Particularly, Deepseek-R1-Zero presents a update sparsity of 86.0%, regardless of directly training from the pretrained base model, namely RL-Zero (DeepSeek-AI et al., 2025), and large-scale training for over 8K steps.
+Particularly, Deepseek-R1-Zero presents a update sparsity of 86.0%, regardless of directly training from the pretrained base model, namely RL-Zero (deepseekai2025deepseekr1incentivizingreasoningcapability), and large-scale training for over 8K steps.
 Although exact training configurations are not always available, we observe that within the same model family, larger models tend to show higher sparsity.
 Importantly, all of these models are trained using full finetuning without any sparsity-promoting regularization techniques or constraints.
-This suggests that the update sparsity is intrinsic to RL training rather than a result of design artifacts.
+This suggests that the update sparsity emerges naturally.
 In contrast, Figure 1 shows that SFT induces dense updates (only 6%-15% sparsity).
-These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities than SFT (Chu et al., 2025; Setlur et al., 2025), possibly by updating substantially fewer parameters.
-Takeaway 1 RL leads to consistently sparse parameter updates (often $>$70% sparsity) while SFT produces dense updates. Sparsity emerges as a characteristic property of RL finetuning in LLMs.
+These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities than SFT (chu2025sftmemorizesrlgeneralizes; setlur2025scalingtesttimecomputeverification), possibly by updating substantially fewer parameters.
+Takeaway 1 RL leads to consistently sparse parameter updates (often >>70% sparsity) while SFT produces dense updates. This sparsity emerges without regularizations or architectural constraints.
 Figure 3: Layerwise and per-parameter-matrix update sparsity for DPO (left) and PRIME (right).
@@
 We next examine how parameter updates in RL are distributed across the model layers and individual parameter matrices (e.g., Q, K, V projections), based on DPO and PRIME models.
-If updates were concentrated in a subset of layers or modules, one could exploit that structure for enhancing the efficiency (Pan et al., 2024).
+If updates were concentrated in a subset of layers or modules, one could exploit that structure for enhancing the efficiency (pan2024lisa).
 Figure 3 shows layerwise and per-matrix sparsity across the models.
@@
 For evaluation, we choose a subset of tasks reported in the original papers for both.
-For DPO we choose the LSAT (Wang et al., 2022), LogiQA (Liu et al., 2021) splits from AGIEval (Zhong et al., 2024), Math split of MMLU Pro (Wang et al., 2024b).
-For PRIME, we report results on the MATH500 (Hendrycks et al., 2021) benchmark across difficulty levels.
+For DPO we choose the LSAT (10.1109/TASLP.2022.3164218), LogiQA (10.5555/3491440.3491941) splits from AGIEval (zhong-etal-2024-agieval), Math split of MMLU Pro (wang2024mmlupro).
+For PRIME, we report results on the MATH500 (hendrycks2021measuring) benchmark across difficulty levels.
 For evaluation in DPO we use olmes.
@@
 They also provide supporting evidence for our Conjecture 1 in §1.
-This finding opens up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (Chen et al., 2022).
+This finding opens up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (chen2022coarseninggranularitystructurallysparse).
 Table 3: Test set performance of $\theta_{\text{full}}$ and $\theta_{\text{sub}}$ trained with DPO and PRIME.
@@
 indicates the difficulty levels of MATH500.
-6cm][t]0.48 Task $\theta_{\text{full}}$ $\theta_{\text{sub}}$ $\Delta$ AGIEval LSAT-AR 21.3 24.8 +3.5 AGIEval LSAT-LR 53.1 54.7 +1.6 AIGEval LogiQA-en 43.5 45.5 +2.0 GPQA 32.8 32.8 +0.0 MMLU Pro Math 50.8 51.6 +0.8 Avg 40.3 41.9 +1.6
+6cm][t]0.48 Task $\theta_{\text{full}}$ $\theta_{\text{sub}}$ $\Delta$ AGIEval LSAT-AR 21.3 24.8 +3.5 AGIEval LSAT-LR 53.1 54.7 +1.6 AGIEval LogiQA-en 43.5 45.5 +2.0 GPQA 32.8 32.8 +0.0 MMLU Pro Math 50.8 51.6 +0.8 Avg 40.3 41.9 +1.6
 (a) DPO
@@
 Unless otherwise mentioned, all ablations were done with a batch size of 32, trained for one epoch with the base model Tulu-3-8B-SFT.
-When varying the training data, we switch between the Tulu preference dataset333allenai/llama-3.1-tulu-3-8b-preference-mixture and the PRIME rollout dataset444 PRIME-RL/Eurus-2-Rollout.
+When varying the training data, we switch between the Tulu preference dataset333allenai/llama-3.1-tulu-3-8b-preference-mixture and the PRIME rollout dataset [^4] PRIME-RL/Eurus-2-Rollout.
 It has model generations for math datasets alongside a label for correctness.
@@
     What factors contribute to the update sparsity observed in RL finetuning?
+[^5] We conjecture that, for the sake of illustration, if one were to perform backpropagation manually on paper with unlimited numerical precision, the resulting parameter updates would be dense.
+In practice, however, modern computers rely on floating-point arithmetic with limited precision.
+As a result, updates with very small magnitudes (e.g., absolute values below $10^{-40}$) cannot be represented and are effectively discarded, hence the empirically observed sparsity of the parameter updates.
+Importantly, such near-zero updates appear to have negligible impact on model performance, as strong results have been achieved without them.
+The equivalent question we address in this section is: What factors contribute to these near-zero parameter updates during RL fine-tuning?
+Figure 5: Update sparsity of intermediate checkpoints of a training run of PRIME. We observe that with more training the sparsity slowly decays.
 We investigate the following factors: gradient clipping, KL regularization towards a reference policy, performing SFT prior to RL, and the number of RL update steps.
@@
 Another important factor is the total number of gradient updates.
-Figure 5: Update sparsity of intermediate checkpoints of a training run of PRIME. We observe that with more training the sparsity slowly decays.
 Gradient clipping and KL regularization.
@@
 Further, SimPO removes the KL term by dropping the base policy normalization in DPO, and as reported in 1 SimPO also produces sparse updates, providing further negative evidence for KL.
+Figure 6: Percentage of updated weights that are outside the final subnetwork across training steps
 Performing SFT before RL.
-A common design choice is to perform SFT on the same data as the subsequent RL Ouyang et al.
-(2022).
+A common design choice is to perform SFT on the same data as the subsequent RL ouyang2022traininglanguagemodelsfollow.
 However, as shown in Table 1, our findings extend to models such as DeepSeek-R1-Zero, which forgoes SFT entirely yet still exhibit high update sparsity.
 This suggests that SFT is not a main contributing factor to the update sparsity in RL finetuning.
-Training duration.
-It is intuitive that with more gradient steps, a model is expected to drift more from the base model.
-Figure 5 shows how update sparsity changes during the training of PRIME.
-As training progresses, update sparsity gradually decreases but eventually shows signs of convergence to around 80%.
-Notably, DeepSeek-R1-Zero undergoes 8K training steps (numbers from Figure 2 in DeepSeek-AI et al.
-(2025)) using GRPO, over 20$\times$ more than PRIME, but shows a comparable update sparsity (86%).
-Therefore, we conjecture that training duration’s impact on update sparsity is more prominent during early training but gradually decreases as training progresses.
-Figure 6: Percentage of updated weights that are outside the final subnetwork across training steps
+[…]
 Figure 6 shows the percentage of updated parameters (relative to model size) that lie outside the final subnetwork.
@@
 While it is possible, though less likely, that all models we study, including DeepSeek-R1-Zero (86.0% update sparsity after 8K steps), are severely undertrained, and that the observed update sparsity would diminish with substantially more training.
-Nonetheless, we question the practicality of this hypothesis since it runs counter to the RL literature arguing against overtraining to prevent overfitting and improve generalization Fu et al.
-(2019).
+Nonetheless, we question the practicality of this hypothesis since it runs counter to the RL literature arguing against overtraining to prevent overfitting and improve generalization pmlr-v97-fu19a.
 Training on in-distribution data.
 Intuitively, when gradients are computed on sequences that the policy already assigns high probabilities to, little update to the parameters would be needed.
-We evaluated two scenarios: (1) rejection sampling, and (2) DPO on out-of-distribution data by not performing SFT pridor to RL.
+We evaluated two scenarios: (1) rejection sampling, and (2) DPO on out-of-distribution data by not performing SFT prior to RL.
 Since we’ve already learned from §3 that DPO with in-distribution data induces sparse updates while SFT (with out-of-distribution data) induces dense ones, this additional experiment can serve as a control group to isolate the factor of training on in- vs.
@@
 Specifically, performing SFT with Qwen/Qwen2.5-Math-7B on rejection sampled in-distribution data yields around 90.0% update sparsity.
-This is reinforced by the examination of a previous work: RAFT++ (Xiong et al., 2024), which performs supervised finetuning with iterative rejection sampling, yields an update sparsity of 69.4%.
+This is reinforced by the examination of a previous work: RAFT++ (xiong2024iterative), which performs supervised finetuning with iterative rejection sampling, yields an update sparsity of 69.4%.
 In contrast, DPO on out-of-distribution data produces dense updates in zephyr-7b-beta models, with a 6.8% update sparsity.
-These findings suggest that training on in-distribution data is a major driver of update sparsity in not only RL, but also SFT.
-Takeaway 4 Training on in-distribution data is a primary reason of update sparsity; KL-divergence regularization and gradient clipping have limited impact.
-Table 5: RFT indicates rejection-sampling fine-tuning (Touvron et al., 2023; Dong et al., 2023), and RAFT++ is iterative RFT A comparative analysis across SFT and DPO as well as in- vs out-of-distribution training shows that in-distribution consistently produces sparse updates.
-Model Method Sparsity (%) SFT/RL In-Dist
+These findings suggest that training on in-distribution data could be a major driver of update sparsity in not only RL, but also SFT.
+Takeaway 4 We conjecture that training on in-distribution data could be a reason of update sparsity; KL-divergence regularization and gradient clipping have limited impact.
+Table 5: RFT indicates rejection‐sampling fine‑tuning (touvron2023llama2openfoundation; dong2023raftrewardrankedfinetuning), and RAFT++ is iterative RFT A comparative analysis across SFT and DPO as well as in- vs out-of-distribution training shows that in-distribution consistently produces sparse updates.
+Model Method Sparsity (%) SFT/RL In‐Dist
 Qwen2.5-Math-7B RFT 91.2 SFT
@@
 Finally, our empirical findings invite a deeper theoretical analysis, with the goal of uncovering theoretical explanations for the update sparsity in RL.
+Lastly, while our observations generally hold, there are confounders (for eg.
+liu2025prorlprolongedreinforcementlearning).
+However the confounding reason is not trivial and is worth exploring as part of future work.
+Further to the best of our observations this phenomenon is largely applicable to all RL settings.
 8 Conclusion
 Our study reveals that RL finetuning in LLMs updates only a sparse subnetwork constituting approximately 5%-30% of total parameters, leaving the rest unchanged.
-This sparsity emerges intrinsically, without explicit spasity promoting techniques such as regularization or structural constraints.
+This sparsity emerges without explicit spasity promoting techniques such as regularization or structural constraints.
 Crucially, finetuning just this subnetwork in isolation reproduces the full model’s performance, aligning closely with original parameter values.
@@
 Our findings highlight that learning from in-distribution samples while training is a key driver of this phenomenon, pointing towards more efficient and effective training strategies in RL-based finetuning of LLMs.
+9 Acknowledgments
+We would like to thank Pavan Jayasinha, Cheng Wang, Abdul Waheed, Shivanshu Shekhar, Alexander Schwing, Gabriel Stanovsky, Roy Schwartz and Gokhan Tur for their valuable feedbacks and suggestions.
+Further we extend our gratitude to the members of ConvAI and Alta lab groups at UIUC for their thoughtful remarks on our draft.
+This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois.
+Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Sections 3-6 provides empirical evidence for the claims made in intro and abstract
+  5. 2.
+    Limitations
+  6. Question: Does the paper discuss the limitations of the work performed by the authors?
+  7. Answer: [Yes]
+  8. Justification: Separate dedicated limiation section included
+  9. 3.
+    Theory assumptions and proofs
+  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  11. Answer: [N/A]
+  12. Justification: The paper doesnt discuss theoretical proofs, most results are empirical
+  13. 4.
+    Experimental result reproducibility
+  14.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  15. Answer: [Yes]
+  16. Justification: All relevant datasets, models are listed, all hyperparameters discussed in detail
+  17. 5.
+    Open access to data and code
+  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  19. Answer: [Yes]
+  20. Justification:
+  21. Guidelines:
+  22. 6.
+    Experimental setting/details
+  23. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  24. Answer: [Yes]
+  25. Justification: All details shared for reproducibility
+  26. Guidelines:
+  27. 7.
+    Experiment statistical significance
+  28. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  29. Answer: [No]
+  30. Justification: Most accuracies are reported on deterministic setups
+  31. 8.
+    Experiments compute resources
+  32. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  33. Answer: [Yes]
+[…]
+  35. 9.
+    Code of ethics
+  36. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  37. Answer: [Yes]
+  38. Justification: Most experiments performed are with publicly released models, datasets and codebases.
+  39. 10.
+    Broader impacts
+  40. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  41. Answer: [N/A]
+  42. Justification: No immediate societal impacts of the work
+  43. 11.
+    Safeguards
+  44.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  45. Answer: [N/A]
+  46. Justification: No such risks
+  47. 12.
+    Licenses for existing assets
+  48.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  49. Answer: [Yes]
+  50. Justification: All works are cited.
+  51. 13.
+    New assets
+  52. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  53. Answer: [N/A]
+  54. Justification: Doesnt release new assets
+  55. 14.
+    Crowdsourcing and research with human subjects
+  56.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  57. Answer: [N/A]
+  58. Justification: Doesn’t involve crowdsourcing.
+  59. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  60.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  61. Answer: [N/A]
+  62. Justification: does not involve crowdsourcing nor research with human subjects
+  63. 16.
+    Declaration of LLM usage
+  64.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  65. Answer: [N/A]
+  66. Justification: Does not involve LLMs usage
 Appendix A Appendix